# Orthogonal Finetuning Made Scalable

### v1

@@ Abstract
Orthogonal finetuning (OFT) offers highly parameter-efficient adaptation while preventing catastrophic forgetting, but its high runtime and memory demands limit practical deployment. We identify the core computational bottleneck in OFT as its weight-centric implementation, which relies on costly matrix-matrix multiplications with cubic complexity. To overcome this, we propose OFTv2, an input-centric reformulation that instead uses matrix-vector multiplications (i.e., matrix-free computation), reducing the computational cost to quadratic. We further introduce the Cayley–Neumann parameterization, an efficient orthogonal parameterization that approximates the matrix inversion in Cayley transform via a truncated Neumann series. These modifications allow OFTv2 to achieve up to 10$\times$ faster training and 3$\times$ lower GPU memory usage without compromising performance. In addition, we extend OFTv2 to support finetuning quantized foundation models and show that it outperforms the popular QLoRA in training stability, efficiency, and memory usage.
@@ 1 Introduction
Another scalability bottleneck in OFT arises from the Cayley parameterization used by Liu et al. (2021a); Qiu et al. (2023); Liu et al. (2024) to preserve orthogonality. While effective, this parameterization involves computing a matrix inverse, which becomes increasingly costly and less numerically stable as weight matrices get larger. To address this, we use a numerically stable yet efficient approximation – the Cayley–Neumann parameterization (CNP) Qiu et al. (2025). By replacing the matrix inverse in the original Cayley transform with a truncated Neumann series, CNP offers improved numerical stability and lower computational cost, particularly in settings where OFT is applied to finetune large foundation models. With CNP, OFTv2 becomes even more scalable and readily applicable for efficient adaptation of such models. In Figure 1, we compare OFT and OFTv2 by performing finetuning tasks on Qwen2.5-7B, which is the largest model the original OFT can finetune within a single Nvidia H100 (80GB). These empirical results demonstrate that OFTv2 achieves substantial GPU memory savings and training speed-up over the original OFT formulation Qiu et al. (2023).
@@
  • Inspired by matrix-free methods that avoid matrix-matrix multiplications in solving linear systems, we propose OFTv2–an input-centric reformulation of OFT that achieves significantly better scalability, with more than 10$\times$ faster training and 3$\times$ lower GPU memory usage.
@@
  • We apply OFTv2 and its quantized variant to different foundation models (including large language models and text-to-image generative models) across various model scale.
@@ 2 Related Work
Orthogonal Finetuning. Qiu et al. (2023); Liu et al. (2024) propose a reparameterization-based method that learns layer-shared orthogonal matrices to transform neurons, yielding strong generalization and stable training. The is motivated by the observation that hyperspherical energy (i.e., a geometric characterization of neurons on the unit sphere) influences generalization Liu et al. (2018, 2021b); Lin et al. (2020); Liu et al. (2023), and that orthogonal transformations keep this energy invariant Liu et al. (2021a). A growing body of research Ma et al. (2024); Yang et al. (2024b); Gorbunov et al. (2024); Yuan et al. (2024); Feng et al. (2025); Raj and Coyle (2025); Lingam et al. (2024); Bini et al. (2024); Su et al. (2024); Liao and Monz (2024) builds upon the core idea of OFT. Figure 2 provides a comparison between OFT and LoRA. OFT achieves parameter efficiency through sparsity, whereas LoRA relies on low-rank structure.
@@ 3.1 Preliminaries
Let $\bm{W}=[\bm{w}_{1},\cdots,\bm{w}_{n}]\in\mathbb{R}^{d\times n}$ be a weight matrix with columns $\bm{w}_{i}\in\mathbb{R}^{d}$. In a linear layer, the forward pass is $\bm{z}=\bm{W}\bm{x}$, where $\bm{x}\in\mathbb{R}^{d}$ is the input and $\bm{z}\in\mathbb{R}^{n}$ is the output. OFT reparameterizes the weight matrix with $\bm{W}_{\text{OFT}}=\bm{R}\bm{W}_{0}$ where $\bm{W}_{0}$ is the pretrained weight matrix and $\bm{R}\in\mathbf{R}^{d\times d}$ is an orthogonal matrix. OFT only learns $\bm{R}$ for adapting the pretrained model to downstream tasks. To enforce orthogonality, Liu et al. (2021b); Qiu et al. (2023); Liu et al. (2024) parameterize $\bm{R}$ using the Cayley transform: $\bm{R}=(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}$, where $\bm{Q}$ is a skew-symmetric matrix satisfying $\bm{Q}=-\bm{Q}^{\top}$. To further improve parameter-efficiency, OFT constrains the orthogonal matrix $\bm{R}$ to have a block-diagonal structure: $\bm{R}=\text{Diag}(\bm{R}_{1},\cdots,\bm{R}_{r})$ where for any $i$, $\bm{R}_{i}\in\mathbb{R}^{b\times b}$ is a small orthogonal matrix and $b\cdot r=d$. Each $\bm{R}_{i}$ can be parameterized using the Cayley transform. This block-diagonal form imposes a sparsity pattern on $\bm{R}$, effectively making it a sparse orthogonal matrix. Leveraging this structure, Liu et al. (2024) further enhance parameter efficiency using butterfly factorization.
@@ 3.2 From Weight-centric Implementation to Input-centric Implementation
$\bm{z}=\underbrace{\overbrace{\bm{W}_{0}^{\top}\bm{R}^{\top}}^{\text{(1) {Weight transform}: matrix-matrix mult.}}\bm{x}}_{\text{(2) {Linear map}: matrix-vector mult.}}$ (1)
@@
$\bm{z}=\underbrace{\bm{W}_{0}^{\top}\overbrace{\bm{R}^{\top}\bm{x}}^{\text{(1){Linear map}: matrix-vector mult.}}}_{\text{(2) {Linear map}: matrix-vector mult.}}$ (2)
where only two matrix-vector multiplications are required, reducing the complexity from cubic to quadratic: $\mathcal{O}(nd+d^{2})$. This simple conceptual shift in implementation entails substantial speed-up in training time and reduction in GPU memory.
@@ 5 Discussions and Intriguing Insights
Figure 4: Results of GPU memory usage for the same finetuning task. (a) OFT, LoRA and OFTv2 on Qwen2.5; (b) QLoRA and QOFT on NF4-quantized Qwen2.5; (c) QLoRA and QOFT on AWQ-qunatized Qwen2.5.
@@ 6.1 GPU Memory Efficiency
As depicted in Figure 1, OFTv2 achieves a 3×3$\times$3 × reduction in GPU memory consumption compared to the original OFT when finetuning the Qwen2.5-7B model. Furthermore, QOFT significantly reduces memory consumption by enabling the orthogonal finetuning of quantized base models. In the following ablation studies comparing against both LoRA and QLoRA baselines, where QLoRA broadly refers to low-rank adaptation of quantized models without being limited to NormalFloat 4-bit quantization, we evaluate the actual GPU memory consumption during finetuning of Qwen2.5 models from 0.5B to 72B parameters. For a comprehensive analysis, we additionally incorporate the widely adopted quantization method AWQ Lin et al. (2024) for activation-aware quantization. The results are summarized in Figure 4. Our experimental results demonstrate that OFTv2 and QOFT achieve memory efficiency comparable to low-rank adaptation methods, with a consistent performance across model scales and data formats.
@@ 6.2 Computational Efficiency
We begin by evaluating the training speed of OFTv2 relative to the original OFT. To this end, we finetune a Qwen2.5-7B model on the OASST1-Guanaco-9K dataset Dettmers et al. (2023) for instruction following and measure the training time. As shown in Figure 1, OFTv2 achieves a 3$\times$ speed-up over the original OFT. We further compare the overall training speed of OFTv2 and LoRA across different model scales and precisions. Settings from both the GSM8K experiment (Table 4) and the OpenR1-Math-220k experiment OpenR1-Team (2025) (Table 5) are used for comparison. Clock times for each setting are reported in Table 1 and Table 2. While low-rank adaptation methods like LoRA benefit from PyTorch’s highly optimized GEMM operations via NVIDIA cuBLAS/cuDNN libraries, the simple designs in OFTv2 significantly narrow this optimization gap in full-precision settings. Notably, OFTv2 outperforms LoRA in quantized settings (Table 2), demonstrating that its quantization-agnostic design effectively leverages underlying quantization-layer optimizations.
@@ 7.1 Encoder-Decoder Model: BART
Qwen2.5-1.5B 8$\times$×H100 01:20:00 01:17:30
Qwen2.5-7B 8$\times$×H100 03:25:00 03:19:30
Qwen2.5-32B 8$\times$×H100 12:51:45 12:27:45
@@
[…]
@@ 7.2 Decoder-only Model: Llama-2 Series
We perform supervised finetuning on the Huggingface OpenR1-Math-220k OpenR1-Team (2025) dataset—a large-scale mathematical reasoning corpus containing challenging problems and two to four reasoning traces distilled from DeepSeek R1 Guo et al. (2025). Following the evaluation protocol of Qwen2.5-Math Yang et al. (2024a), we report pass@1 performance on established math benchmarks: CMATH Wei et al. (2023), AMC23 Project-Numina , AQUA Ling et al. (2017), Olympiad Bench He et al. (2024), Gaokao 2023 En Liao et al. (2024), and Minerva Math Lewkowycz et al. (2022). Finetuning was only performed on NormalFloat 4-bit quantized base models due to the substantial memory requirements imposed by the large context window size (16384), necessary for training on a reasoning dataset. The results are reported in Table 5. The baseline method refers to the pre-trained Qwen2.5 models without any continual training. We observe that QOFT consistently outperforms both QLoRA and the base model across all evaluated scales and tasks, despite using significantly fewer trainable parameters. For instance, on the Qwen2.5-7B instruction-tuned model, QOFT achieves a 96.9% SAT Math accuracy compared to QLoRA’s 68.8%, while utilizing only 17.55M parameters (57% fewer than QLoRA’s 40.37M). This advantage scales robustly: the Qwen2.5-32B variant finetuned with QOFT attains 100% SAT Math accuracy, surpassing both the baseline (65.6%) and QLoRA (96.9%). These gains persist across mathematical reasoning tasks (e.g., 70.0% on AMC23 for QOFT-32B vs. QLoRA’s 40.0%), suggesting that orthogonal adaptation in quantized space better preserves the model’s reasoning capabilities compared to low-rank adaptation. The results demonstrate QOFT’s dual strength: parameter efficiency without sacrificing task performance, particularly in the quantized setting. In contrast, QLoRA-finetuned models can exhibit training instabilities Li et al. (2023), leading to model collapse where their performance fell below the base model. Appendix B gives more results on finetuning math-specific Qwen2.5 models.
@@
To assay the generality of the proposed methods across modalities, we perform Dreambooth Ruiz et al. (2023) finetuning on the latest Stable Diffusion 3.5 models Esser et al. (2024). Dreambooth finetunes text-to-image models using a limited set of images depicting the same subject. This process binds the subject to a unique token identifier, enabling subject-driven generation where the model synthesizes this subject in novel scenes beyond the training data. Qualitative results are shown in Figure 5 and Appendix C. We also report the actual peak GPU memory usage during the finetuning process in Appendix C. For finetuning the NormalFloat 4-bit quantized Stable Diffusion 3.5 Large model, QOFT requires slightly less GPU memory ($35.02$ GB) than the QLoRA method ($35.03$ GiB).
@@ Appendix A Experimental Details
We employed the Adam optimizer Kingma and Ba (2015) for all our training runs. The specific hyperparameters used for each experiment are detailed in the tables referenced below. These include learning rates, batch sizes, number of training epochs, and method-specific configurations: the rank $r$ for LoRA-based methods and the block size $b$ for OFTv2/QOFT. If not explicitly specified, the $r$ for LoRA-based methods is 16 and the block size $b$ for OFTv2/QOFT is set as 32. For the Wikitext dataset, hyperparameters are listed in Table 8. For the GSM8K dataset, hyperparameters are listed in Table 9. For the XSum dataset, hyperparameters are listed in Table 6. For the CNN/DailyMail dataset, hyperparameters are listed in Table 7. Since it is known that merging QLoRA adapter weights to its quantized base models leads to performance degradation [^1] Comparison of merging methods: https://kaitchup.substack.com/p/lora-adapters-when-a-naive-merge and distort the real performance, for every experiment, we evaluate the fine-tuned model without merging the trainable parameters, but load them as extra adapter layers.
@@
Appendix B Mathematical Reasoning with Qwen2.5
@@
Appendix C Subject-driven Generation with Stable diffusion 3.5

### v2

@@ Abstract
[…]
@@ 1 Introduction
Another scalability bottleneck in OFT arises from the Cayley parameterization used by Liu et al. (2021a); Qiu et al. (2023); Liu et al. (2024) to preserve orthogonality. While effective, this parameterization involves computing a matrix inverse, which becomes increasingly costly and less numerically stable as weight matrices get larger. To address this, we use a numerically stable yet efficient approximation – the Cayley–Neumann parameterization (CNP) Qiu et al. (2025). By replacing the matrix inverse in the original Cayley transform with a truncated Neumann series, CNP offers improved numerical stability and lower computational cost, particularly in settings where OFT is applied to finetune large foundation models. With CNP, OFTv2 becomes even more scalable and readily applicable for efficient adaptation of such models. In Figure 1, we compare OFT and OFTv2 by performing finetuning tasks on Qwen2.5-7B, which is the largest model that the original OFT can finetune within a single Nvidia H100 (80GB). These empirical results demonstrate that OFTv2 achieves substantial GPU memory savings and training speed-up over the original OFT formulation Qiu et al. (2023).
@@
[…]
@@
  • We apply OFTv2 and its quantized variant to different foundation models (including large language models and text-to-image generative models) across various model scales.
@@ 2 Related Work
Orthogonal Finetuning. Qiu et al. (2023); Liu et al. (2024) propose a reparameterization-based method that learns layer-shared orthogonal matrices to transform neurons, yielding strong generalization and stable training. The is motivated by the observation that hyperspherical energy (i.e., a geometric characterization of neurons on the unit sphere) influences generalization Liu et al. (2018, 2021b); Lin et al. (2020); Liu et al. (2023), and that orthogonal transformations keep this energy invariant Liu et al. (2021a). A growing body of research Ma et al. (2024); Yang et al. (2024b); Gorbunov et al. (2024); Yuan et al. (2024); Feng et al. (2025); Raj and Coyle (2025); Lingam et al. (2024); Bini et al. (2024); Su et al. (2024); Liao and Monz (2024) builds upon the core idea of OFT. Figure 2 provides a comparison between OFT and LoRA. OFT achieves parameter efficiency through sparsity, whereas LoRA relies on a low-rank structure.
@@ 3.1 Preliminaries
Let $\bm{W}=[\bm{w}_{1},\cdots,\bm{w}_{n}]\in\mathbb{R}^{d\times n}$ be a weight matrix with columns $\bm{w}_{i}\in\mathbb{R}^{d}$. In a linear layer, the forward pass is $\bm{z}=\bm{W}\bm{x}$, where $\bm{x}\in\mathbb{R}^{d}$ is the input and $\bm{z}\in\mathbb{R}^{n}$ is the output. OFT reparameterizes the weight matrix with $\bm{W}_{\text{OFT}}=\bm{R}\bm{W}_{0}$ where $\bm{W}_{0}$ is the pretrained weight matrix and $\bm{R}\in\mathbf{R}^{d\times d}$ is an orthogonal matrix. OFT only learns $\bm{R}$ for adapting the pretrained model to downstream tasks. To enforce orthogonality, Liu et al. (2021b); Qiu et al. (2023); Liu et al. (2024) parameterize $\bm{R}$ using the Cayley transform: $\bm{R}=(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}$, where $\bm{Q}$ is a skew-symmetric matrix satisfying $\bm{Q}=-\bm{Q}^{\top}$. To further improve parameter-efficiency, OFT constrains the orthogonal matrix $\bm{R}$ to have a block-diagonal structure: $\bm{R}=\text{Diag}(\bm{R}_{1},\cdots,\bm{R}_{r})$ where for any $i$, $\bm{R}_{i}\in\mathbb{R}^{b\times b}$ is a small orthogonal matrix and $b\cdot r=d$. Each $\bm{R}_{i}$ can be parameterized using the Cayley transform. This block-diagonal form imposes a sparsity pattern on $\bm{R}$, effectively making it a sparse orthogonal matrix. Leveraging this structure, Liu et al. (2024) further enhances parameter efficiency using butterfly factorization.
@@ 3.2 From Weight-centric Implementation to Input-centric Implementation
[…]
@@
[…]
where only two matrix-vector multiplications are required, reducing the complexity from cubic to quadratic: $\mathcal{O}(nd+d^{2})$. This simple conceptual shift in implementation entails a substantial speed-up in training time and reduction in GPU memory.
@@ 5 Discussions and Intriguing Insights
Figure 4: Results of GPU memory usage for the same finetuning task. (a) OFT, LoRA and OFTv2 on Qwen2.5; (b) QLoRA and QOFT on NF4-quantized Qwen2.5; (c) QLoRA and QOFT on AWQ-quantized Qwen2.5.
@@ 6.1 GPU Memory Efficiency
As depicted in Figure 1, OFTv2 achieves a $3\times$ reduction in GPU memory consumption compared to the original OFT when finetuning the Qwen2.5-7B model. Furthermore, QOFT significantly reduces memory consumption by enabling the orthogonal finetuning of quantized base models. In the following ablation studies comparing against both LoRA and QLoRA baselines, where QLoRA broadly refers to low-rank adaptation of quantized models without being limited to NormalFloat 4-bit quantization, we evaluate the actual GPU memory consumption during finetuning of Qwen2.5 models from 0.5B to 72B parameters. For a comprehensive analysis, we additionally incorporate the widely adopted quantization method AWQ Lin et al. (2024) for activation-aware quantization. The results are summarized in Figure 4. Our experimental results demonstrate that OFTv2 and QOFT achieve memory efficiency comparable to low-rank adaptation methods, with a consistent performance across model scales and data formats.
@@ 6.2 Computational Efficiency
[…]
@@ 7.1 Encoder-Decoder Model: BART
[…]
@@
# $uparrow # Params XSum\uparrow CNN/DailyMail\uparrow$
@@ 7.2 Decoder-only Model: Llama-2 Series
We perform supervised finetuning on the Huggingface OpenR1-Math-220k OpenR1-Team (2025) dataset—a large-scale mathematical reasoning corpus containing challenging problems and two to four reasoning traces distilled from DeepSeek R1 Guo et al. (2025). Following the evaluation protocol of Qwen2.5-Math Yang et al. (2024a), we report pass@1 performance on established math benchmarks: CMATH Wei et al. (2023), AMC23 Project-Numina , AQUA Ling et al. (2017), Olympiad Bench He et al. (2024), Gaokao 2023 En Liao et al. (2024), and Minerva Math Lewkowycz et al. (2022). Finetuning was only performed on NormalFloat 4-bit quantized base models due to the substantial memory requirements imposed by the large context window size (16384), necessary for training on a reasoning dataset. The results are reported in Table 5. The baseline method refers to the pre-trained Qwen2.5 models without any continual training. We observe that QOFT consistently outperforms both QLoRA and the base model across all evaluated scales and tasks, despite using significantly fewer trainable parameters. For instance, on the Qwen2.5-7B instruction-tuned model, QOFT achieves a 96.9% SAT Math accuracy compared to QLoRA’s 68.8%, while utilizing only 17.55M parameters (57% fewer than QLoRA’s 40.37M). This advantage scales robustly: the Qwen2.5-32B variant finetuned with QOFT attains 100% SAT Math accuracy, surpassing both the baseline (65.6%) and QLoRA (96.9%). These gains persist across mathematical reasoning tasks (e.g., 70.0% on AMC23 for QOFT-32B vs. QLoRA’s 40.0%), suggesting that orthogonal adaptation in quantized space better preserves the model’s reasoning capabilities compared to low-rank adaptation. The results demonstrate QOFT’s dual strength: parameter efficiency without sacrificing task performance, particularly in the quantized setting. In contrast, QLoRA-finetuned models can exhibit training instabilities Li et al. (2023), leading to model collapse where their performance fell below the base model. Appendix C gives more results on finetuning math-specific Qwen2.5 models.
@@
To assay the generality of the proposed methods across modalities, we perform Dreambooth Ruiz et al. (2023) finetuning on the latest Stable Diffusion 3.5 models Esser et al. (2024). Dreambooth finetunes text-to-image models using a limited set of images depicting the same subject. This process binds the subject to a unique token identifier, enabling subject-driven generation where the model synthesizes this subject in novel scenes beyond the training data. Qualitative results are shown in Figure 5 and Appendix D. We also report the actual peak GPU memory usage during the finetuning process in Appendix D. For finetuning the NormalFloat 4-bit quantized Stable Diffusion 3.5 Large model, QOFT requires slightly less GPU memory ($35.02$ GB) than the QLoRA method ($35.03$ GiB).
@@ Appendix A Experimental Details
We employed the Adam optimizer Kingma and Ba (2015) for all our training runs. The specific hyperparameters used for each experiment are detailed in the tables referenced below. These include learning rates, batch sizes, number of training epochs, and method-specific configurations: the rank $r$ for LoRA-based methods and the block size $b$ for OFTv2/QOFT. If not explicitly specified, the $r$ for LoRA-based methods is 16 and the block size $b$ for OFTv2/QOFT is set as 32. For the Wikitext dataset, hyperparameters are listed in Table 8. For the GSM8K dataset, hyperparameters are listed in Table 9. For the XSum dataset, hyperparameters are listed in Table 6. For the CNN/DailyMail dataset, hyperparameters are listed in Table 7. Since it is known that merging QLoRA adapter weights to its quantized base models leads to performance degradation [^1] Comparison of merging methods: https://kaitchup.substack.com/p/lora-adapters-when-a-naive-merge and distorts the real performance, for every experiment, we evaluate the fine-tuned model without merging the trainable parameters, but load them as extra adapter layers.
@@
Appendix B Effect of Neumann Series Terms in Orthogonal Parameterization
OFTv2 employs the Cayley-Neumann parameterization to improve the training efficiency; the number of Neumann series terms becomes a hyperparameter. We conducted an additional ablation study to evaluate the impact of the number of Neumann series terms on finetuning performance for WikiText. The results are reported in Table 10. We observe that when the number of Neumann terms is too small (e.g., 2), the approximation error to orthogonality slightly degrades performance. For the experiments reported in the main paper, we used five Neumann terms, which we found to be well-suited across all evaluated tasks.
Model Method 2 terms 3 terms 4 terms 5 terms 6 terms
Llama 2 7B OFTv2 6.22 6.15 6.14 6.13 6.14
Llama 2 13B OFTv2 5.11 5.00 4.99 4.98 4.99
Llama 2 7B QOFT 5.70 5.62 5.58 5.60 5.61
Llama 2 13B QOFT 5.14 5.02 5.04 5.05 5.05
Table 10: Effect of Neumann Series Terms on the Llama-2 Models
Appendix C Mathematical Reasoning with Qwen2.5
@@ Appendix B Effect of Neumann Series Terms in Orthogonal Parameterization
Appendix D Subject-driven Generation with Stable diffusion 3.5